6/(x^2-9-1/(x-3)=0
Answer by mananth(16949)

6/(x^2-9) - 1/(x-3)=0
6/(x+3)(x-3) - 1/(x-3)=0
LCM (x+3)(x-3)
.
6-x-3 /(x+3)(x-3)=0
-x+3 / (x+3)(x-3)=0
-1(x-3)/ (x+3)(x-3)=0
-1/(x-3)=0
x-3=-1
x=2